OpenStudy (0241595):

Help!! Prove the Converse of the Pythagorean Theorem using similar triangles. The Converse of the Pythagorean Theorem states that when the sum of the squares of the lengths of the legs of the triangle equals the squared length of the hypotenuse, the triangle is a right triangle. Be sure to create and name the appropriate geometric figures.

OpenStudy (welshfella):

I can prove it using congruent triangles.

OpenStudy (0241595):

That's fine, what is it?

OpenStudy (welshfella):



OpenStudy (0241595):

So what's the proof?

OpenStudy (welshfella):

let ABC be a triangle where AB^2 = CB^2 + AC^2 WE say that angle ACB is right angled Draw a line segment BCD so that BC = CD Now join AD

OpenStudy (welshfella):



OpenStudy (welshfella):

Since DC = CB the square on BC must equal the square on DC . DC^2 = CB^2. Now Add the square on AC to each then DA^2 + DC^2 = DA^2 + CB^2

OpenStudy (0241595):

Wait but if you're adding the square of AC then how do you get DA^2 + DC^2 = DA^2 + CB^2?

OpenStudy (welshfella):

SOrry the last statement should read AC^2 + DC^2 = AC^2 + CB^2

OpenStudy (welshfella):

But we are given that AB^2 = AC^2 + CB^2 and therefore AB^2 = AD^2 ( I think you call that the transitive property of equality in the US)

OpenStudy (welshfella):

so AD = AB and we are given that DC=BC and AD = AD therefore the 2 triangle are congruent by SSS

OpenStudy (0241595):

Ohh ok, thanks that helped a lot :)

OpenStudy (welshfella):

Finally the 2 angles ACD = ACB so they both must be right angles therefore triangle ABC is a right angle.

OpenStudy (0241595):

Yeah thanks I was mostly confused about one of the parts but you explained it great, thank you

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

yeah I had a mental 'blip'.

OpenStudy (welshfella):

- or at my age its called a 'senior moment' LOL

OpenStudy (welshfella):

THis proof is due to Euclid .